OpenStudy (anonymous):

Solve the exponential equations symbolically. Give exact values as answers. (a) 2^x = 5 (b) 7e^(2x) - 8 = 377

OpenStudy (anonymous):

2^x = 5 ln(2^x) = ln(5) x*ln(2) = ln(5) x = ln(5)/ln(2)

OpenStudy (anonymous):

ahh ok I think I remember

OpenStudy (anonymous):

B. is killing me though.

OpenStudy (anonymous):

7e^(2x) - 8 = 377 7e^(2x) = 377+8 e^(2x) = 385/7 e^(2x) = 55 ln(e^(2x)) = ln(55) //ln(e^a) = a 2x = ln(55) x = (ln(55))/2

OpenStudy (anonymous):

ahh ok so I have to get everything off first then I do ln. How in the world didn't I see that.